OpenStudy (anonymous):

sqrt(x) - sqrt(3x-3) = 1

OpenStudy (anonymous):

\[\sqrt{x}-\sqrt{3x-3}=1\] \[\sqrt{x}=1+\sqrt{3x-3}\] Now raise both sides to the power of 2 to get rid of the square root.